OpenStudy (sohailiftikhar):

Evaluate the following integral exactly, no approximation, and justify lower limit is zero and uper limit is infinity ∫e−x3√x(1−sinx)(1+2sinx−cos2x) dx

OpenStudy (sohailiftikhar):

so dude can you solve it huh ? :)

imqwerty (imqwerty):

is it this- \[\int\limits_{0}^{\infty}[e-x \sqrt[3]{x(1-sinx)(1+2sinx-\cos2x)}]dx\]

OpenStudy (sohailiftikhar):